# MVG-CRPS: A Robust Loss Function for Multivariate Probabilistic Forecasting

### v1

@@ 5 CONCLUSION
In this paper, we have introduced MVG-CRPS, a robust loss function specifically designed for multivariate Gaussian outputs in probabilistic forecasting. Addressing the high sensitivity of the log-score to outliers and the inefficiency of the energy score, our approach leverages the CRPS to enhance robustness without compromising computational efficiency. The closed-form expression of MVG-CRPS based on neural network outputs facilitates its seamless integration into deep learning models. Our experimental results on real-world datasets demonstrate that MVG-CRPS outperforms existing loss functions like log-score and energy score, offering improved accuracy and robustness in probabilistic forecasting tasks. While we have focused on forecasting applications, the generality of our method allows it to be extended to a wide range of probabilistic regression problems involving multivariate Gaussian distributed responses.
For future work, we plan to explore the use of copulas to transform continuous distributions into Gaussian ones, thereby broadening the applicability of our approach to non-Gaussian settings. Additionally, we aim to investigate better parameterizations for covariance matrices to further enhance model performance and scalability. Overall, MVG-CRPS provides a general solution for multivariate Gaussian distribution forecasting, contributing to the development of more robust and accurate probabilistic models in the presence of outliers.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [No]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Not Applicable]
    2. (b)
      Complete proofs of all theoretical results. [Not Applicable]
    3. (c)
      Clear explanations of any assumptions. [Not Applicable]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]

### v2

@@ 5 Conclusion
In this paper, we introduced MVG-CRPS, a loss function designed for multivariate Gaussian outputs in probabilistic forecasting. By addressing the sensitivity of the log-score to outliers and the inefficiency of the energy score, we leveraged the CRPS to improve robustness without sacrificing computational efficiency. Its closed-form expression based on neural network outputs allows seamless integration into deep learning models. Our experiments on real-world datasets show that MVG-CRPS provides improved accuracy and robustness in probabilistic forecasting tasks.
Furthermore, MVG-CRPS can be used for general tasks, including robust Gaussian process regression, by replacing the negative marginalized likelihood. While our focus has been on forecasting applications, the general nature of MVG-CRPS makes it applicable to a wide range of probabilistic regression problems involving multivariate Gaussian responses.
For future work, we plan to employ copulas to transform continuous distributions into Gaussian ones, extending our approach to non-Gaussian settings. We also aim to investigate improved parameterizations for covariance matrices to further enhance scalability. A limitation of our current approach is scalability with large batch sizes. Although a low-rank-plus-diagonal structure can be leveraged for the log-score through the Woodbury identity and the matrix-determinant lemma (when the batch size exceeds the factor dimension), MVG-CRPS does not benefit from this structure due to the SVD operation. A possible solution is to adopt an isotropic noise parameterization, i.e., $\boldsymbol{\Sigma}=\boldsymbol{L}\boldsymbol{L}^{\top}+\sigma^{2}\mathbf{I}$, which enables more efficient computation of the SVD.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
A Properties of MVG-CRPS A
B Dataset Details B
C Experiment Details C
C.1 Benchmark Models C.1
C.2 Hyperparameters C.2
C.3 Training Procedure C.3
C.4 Covariance Parameterization C.4
C.5 SVD and Gradient Calculation C.5
C.6 Naive Baseline Description C.6
D Additional Experiments D
Appendix A Properties of MVG-CRPS
Theorem A.1.
Let $\mathbf{z}\sim\mathcal{N}\left(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p}\right)$ be a true $N$-variate Gaussian distribution where the covariance admits eigen-decomposition $\boldsymbol{\Sigma}_{p}=\boldsymbol{U}_{p}\boldsymbol{S}_{p}\boldsymbol{U}_{p}^{\top}$, with $\boldsymbol{S}_{p}=\operatorname{diag}\left(\boldsymbol{\lambda}_{p}\right)$ containing nonincreasing eigenvalues $\boldsymbol{\lambda}_{p}=\left[\lambda_{1}^{p},\ldots,\lambda_{N}^{p}\right]^{\top}$ and $\boldsymbol{U}_{p}$ being the corresponding orthonormal matrix. Consider a predictive Gaussian distribution $\mathcal{N}\left(\boldsymbol{\mu}_{q},\boldsymbol{\Sigma}_{q}\right)$, where covariance $\boldsymbol{\Sigma}_{q}$ admits the eigen-decomposition $\boldsymbol{\Sigma}_{q}=\boldsymbol{U}_{q}\boldsymbol{S}_{q}\boldsymbol{U}_{q}^{\top}$ with $\boldsymbol{S}_{q}=\operatorname{diag}\left(\boldsymbol{\lambda}_{q}\right)$. Define the transformed variable $\mathbf{v}=\boldsymbol{U}_{q}^{\top}\left(\mathbf{z}-\boldsymbol{\mu}_{q}\right)=\left[v_{1},\ldots,v_{N}\right]^{\top}$. The proposed MVG-CRPS
$\mathcal{L}\left(\mathcal{N}\left(\boldsymbol{\mu}_{q},\boldsymbol{\Sigma}_{q}\right),\mathbf{z}\right)=\sum_{i=1}^{N}\operatorname{CRPS}\left(\Phi\left(0,\lambda_{i}^{q}\right),v_{i}\right)$
is proper and strictly proper for multivariate Gaussian distributions.
Proof.
Since $\mathbf{z}\sim\mathcal{N}\left(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p}\right)$, we have the transformed variable $\mathbf{v}\sim\mathcal{N}\left(\boldsymbol{\mu}_{v},\boldsymbol{\Sigma}_{v}\right)$ with $\boldsymbol{\mu}_{v}=\boldsymbol{U}_{q}^{\top}(\boldsymbol{\mu}_{p}-\boldsymbol{\mu}_{q})=\left[\nu_{1},\ldots,\nu_{N}\right]^{\top}$ and $\boldsymbol{\Sigma}_{v}=\boldsymbol{U}_{q}^{\top}\boldsymbol{\Sigma}_{p}\boldsymbol{U}_{q}=\boldsymbol{U}_{q}^{\top}\boldsymbol{U}_{p}\boldsymbol{S}_{p}\boldsymbol{U}_{p}^{\top}\boldsymbol{U}_{q}=\boldsymbol{U}_{v}\boldsymbol{S}_{p}\boldsymbol{U}_{v}^{\top}$, where $\boldsymbol{U}_{v}=\boldsymbol{U}_{q}^{\top}\boldsymbol{U}_{p}$ is an orthonormal matrix. Thus, each $v_{i}$ has a marginal distribution $v_{i}\sim\mathcal{N}\left(\nu_{i},\phi_{i}\right)$ with $\boldsymbol{\phi}=\operatorname{diag}(\boldsymbol{\Sigma}_{v})=\operatorname{diag}(\boldsymbol{U}_{v}\boldsymbol{S}_{p}\boldsymbol{U}_{v}^{\top})=\left[\phi_{1},\ldots,\phi_{N}\right]^{\top}$ for $i=1,\ldots,N$. Taking the expectation of $\mathcal{L}\left(\mathcal{N}\left(\boldsymbol{\mu}_{q},\boldsymbol{\Sigma}_{q}\right),\mathbf{z}\right)$ under the true distribution, we have
$\begin{split}\mathbb{E}_{\mathbf{z}\sim\mathcal{N}\left(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p}\right)}\left[\mathcal{L}\left(\mathcal{N}\left(\boldsymbol{\mu}_{q},\boldsymbol{\Sigma}_{q}\right),\mathbf{z}\right)\right]&=\sum_{i=1}^{N}\mathbb{E}_{v_{i}\sim\mathcal{N}\left(\nu_{i},\phi_{i}\right)}\left[\operatorname{CRPS}\left(\Phi\left(0,\lambda_{i}^{q}\right),v_{i}\right)\right]\\ &\geq\sum_{i=1}^{N}\mathbb{E}_{v_{i}\sim\mathcal{N}\left(\nu_{i},\phi_{i}\right)}\left[\operatorname{CRPS}\left(\Phi\left(\nu_{i},\phi_{i}\right),v_{i}\right)\right]\\ &=\sum_{i=1}^{N}\mathbb{E}_{\eta_{i}\sim\mathcal{N}\left(0,\phi_{i}\right)}\left[\operatorname{CRPS}\left(\Phi\left(0,\phi_{i}\right),\eta_{i}\right)\right]\\ &=\mathbb{E}_{v\sim\mathcal{N}\left(0,1\right)}\left[\operatorname{CRPS}\left(\Phi,v\right)\right]\times\sum_{i=1}^{N}\sqrt{\phi_{i}}\\ &\geq\mathbb{E}_{v\sim\mathcal{N}\left(0,1\right)}\left[\operatorname{CRPS}\left(\Phi,v\right)\right]\times\sum_{i=1}^{N}\sqrt{\lambda_{i}^{p}}\\ &=\mathbb{E}_{\mathbf{z}\sim\mathcal{N}\left(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p}\right)}\left[\mathcal{L}\left(\mathcal{N}\left(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p}\right),\mathbf{z}\right)\right].\end{split}$ (18)
The first inequality is a direct result of CRPS being a strictly proper scoring rule for univariate Gaussian distributions. We now prove the second inequality.
Recall that $\boldsymbol{\phi}=\operatorname{diag}(\boldsymbol{\Sigma}_{v})$ and $\boldsymbol{\Sigma}_{v}=\boldsymbol{U}_{q}^{\top}\boldsymbol{\Sigma}_{p}\boldsymbol{U}_{q}$. Let $\boldsymbol{\phi}^{*}$ be the monotone nonincreasing rearrangement of $\boldsymbol{\phi}$. By the Schur-Horn theorem Horn (1954), the vector $\boldsymbol{\phi}^{*}$ is submajorized by the eigenvalues $\boldsymbol{\lambda}_{p}$:
$\sum_{i=1}^{k}\phi_{i}^{*}\leq\sum_{i=1}^{k}\lambda_{i}^{p},$
for $k=1,2,\dots,N-1$, and
$\sum_{i=1}^{N}\phi_{i}^{*}=\sum_{i=1}^{N}\lambda_{i}^{p}.$
Since $f(x)=\sqrt{x}$ is a concave function, Karamata’s majorization inequality yields
$\sum_{i=1}^{N}\sqrt{\lambda_{i}^{p}}\leq\sum_{i=1}^{N}\sqrt{\phi_{i}^{*}}=\sum_{i=1}^{N}\sqrt{\phi_{i}},$ (19)
which proves the second inequality in Eq. (18).
Equality in Eq. (19) is obtained if, for every $i$, $\phi_{i}^{*}=\lambda_{i}^{p}$. By the Schur-Horn theorem, this forces $\boldsymbol{\Sigma}_{v}$ has to be a diagonal matrix (Theorem 4.3.45 in Horn & Johnson (2012)). Meanwhile, the CRPS inequality in Eq. (18) is tight exactly when, for every $i$, $\nu_{i}=0$ and $\phi_{i}=\lambda_{i}^{q}$, implying that $\boldsymbol{U}_{q}^{\top}(\boldsymbol{\mu}_{p}-\boldsymbol{\mu}_{q})=\boldsymbol{0}$ and $\operatorname{diag}(\boldsymbol{\Sigma}_{v})=\operatorname{diag}(\boldsymbol{S_{q}})$. Since $\boldsymbol{\Sigma}_{v}$ is diagonal, we have $\boldsymbol{\Sigma}_{v}=\boldsymbol{U}_{q}^{\top}\boldsymbol{\Sigma}_{p}\boldsymbol{U}_{q}=\boldsymbol{S}_{q}$, hence $\boldsymbol{\Sigma}_{p}=\boldsymbol{\Sigma}_{q}$. Therefore, all equalities hold if and only if $\boldsymbol{\mu}_{p}=\boldsymbol{\mu}_{q}$ and $\boldsymbol{\Sigma}_{p}=\boldsymbol{\Sigma}_{q}$. This confirms that the proposed scoring rule is proper and strictly proper for multivariate Gaussian distributions. ∎
Appendix B Dataset Details
We conducted experiments on a diverse collection of real-world datasets sourced from GluonTS (Alexandrov et al., 2020). These datasets are commonly used for benchmarking time series forecasting models, following their default configurations in GluonTS, which include granularity, prediction horizon ($Q$), and the number of rolling evaluations. For each dataset, we sequentially split the data into training, validation, and testing sets, ensuring that the temporal length of the validation set matched that of the testing set. The temporal length of the testing set was based on the prediction horizon and the required number of rolling evaluations. For example, the testing horizon for the $\mathtt{traffic}$ dataset is calculated as $24+7-1=30$ time steps. Consequently, the model generates 24-step predictions ($Q$) sequentially, with 7 distinct consecutive prediction start points, corresponding to 7 forecast instances. In our experiments, we aligned the conditioning range ($P$) with the prediction horizon ($Q$), consistent with the default setting in GluonTS (i.e., $P=Q$). Each dataset was normalized using the mean and standard deviation computed from each time series in the training set, and predictions were rescaled to their original values for evaluation. Table 4 summarizes the statistics of all datasets.
Table 4: Dataset summary.
Dataset Granularity # of time series # of time steps $Q$ Rolling evaluation
$\mathtt{elec\_au}$ 30min 5 232,272 60 56
$\mathtt{cif\_2016}$ monthly 72 120 12 1
$\mathtt{electricity}$ hourly 370 5,857 24 7
$\mathtt{elec\_weekly}$ weekly 321 156 8 3
$\mathtt{exchange\_rate}$ workday 8 6,101 30 5
$\mathtt{kdd\_cup}$ hourly 270 10,920 48 7
$\mathtt{m1\_yearly}$ yearly 181 169 6 1
$\mathtt{m3\_yearly}$ yearly 645 191 6 1
$\mathtt{nn5\_daily}$ daily 111 791 56 5
$\mathtt{saugeenday}$ daily 1 23,741 30 5
$\mathtt{sunspot}$ daily 1 73,924 30 5
$\mathtt{tourism}$ quarterly 427 131 8 1
$\mathtt{traffic}$ hourly 963 4,025 24 7
$\mathtt{covid}$ daily 266 212 30 5
$\mathtt{elec\_hourly}$ hourly 321 26,304 48 7
$\mathtt{m4\_hourly}$ hourly 414 1,008 48 7
$\mathtt{pedestrian}$ hourly 66 96,432 48 7
$\mathtt{taxi\_30min}$ 30min 1214 1,637 24 56
$\mathtt{uber\_hourly}$ hourly 262 8,343 24 7
$\mathtt{wiki}$ daily 2000 792 30 5
Appendix C Experiment Details
C.1 Benchmark Models
The input to benchmark models includes lagged time series values and covariates that encode time and series identification. The number of lagged values is determined by the granularity of each dataset. Specifically, we use lags of [1, 24, 168] for hourly data, [1, 7, 14] for daily data, and [1, 2, 4, 12, 24, 48] for data with sub-hourly granularity. For all other datasets, only lag-1 values are used.
For datasets with hourly or finer granularity, we include the hour of the day and day of the week. For daily datasets, only the day of the week is used. Each time series is uniquely identified by a numeric identifier. All features are encoded as single values; for example, the hour of the day takes values between [0, 23]. These features are concatenated with the model input at each time step to form the model input vector $\mathbf{y}_{t}$ (Salinas et al., 2019; Zheng & Sun, 2024).
Our method requires a state vector $\mathbf{h}_{i,t}$ to generate the parameters for the predictive distribution. To achieve this, we employ different neural architectures: RNNs and Transformer decoders, both of which maintain autoregressive properties for the multivariate autoregressive forecasting task, and MLPs for the univariate Seq2Seq forecasting task. Specifically, we use the GPVar model (Salinas et al., 2019) as our RNN benchmark, the GPT model (Radford et al., 2018) for the decoder-only Transformer, and the N-HiTS model (Challu et al., 2023) for the MLPs. All models are trained to output $\mathbf{h}_{i,t}$, which is used to parameterize the predictive distribution.
C.2 Hyperparameters
All model parameters are optimized using the Adam optimizer with $l_{2}$ regularization set to 1e-8, and gradient clipping applied at 10.0. For all methods, we cap the total number of gradient updates at 10,000 and reduce the learning rate by a factor of 2 after 500 consecutive updates without improvement. Table 5 provides the hyperparameter values that remain fixed across all datasets. We did NOT tune the hyperparameters specifically to favor the proposed loss. Instead, we used the same hyperparameters as those in GPVar (Salinas et al., 2019), which were originally tuned for the log-score. Keeping the hyperparameters consistent across losses ensures that any observed improvements are attributable to the loss function itself, rather than differences in hyperparameter settings.
Table 5: Hyperparameters values.
Hyperparameter Value
learning rate 1e-3
hidden size 40
$\mathtt{n\_layers}$ (RNN/Transformer decoder/MLP) 2
$\mathtt{n\_heads}$ (Transformer) 2
rank ($R$) 10
sampling dimension ($B$) 20
dropout 0.01
batch size 16
C.3 Training Procedure
Compute Resources All models were trained in an Anaconda environment using one AMD Ryzen Threadripper PRO 5955WX CPU and four NVIDIA RTX A5000 GPUs, each with 24 GB of memory.
Batch Size Following the method used in GPVar (Salinas et al., 2019), we set the sample slice size to $B=20$ time series and used a batch size of 16. Since our data sampler processes one slice of time series at a time rather than sampling 16 slices simultaneously, we set $\mathtt{accumulate\_grad\_batches}$ to 16, effectively achieving a batch size of 16.
Training Loop During each epoch, the model is trained on up to 400 batches from the training set, followed by the computation of the $\mathtt{valid\_loss}$ on the validation set. Training is halted when one of the following conditions is met:
  • A total of 10,000 gradient updates has been reached,
  • No improvement in the validation set $\mathtt{valid\_loss}$ is observed for 10 consecutive epochs.
The final model is the one that achieves the lowest $\mathtt{valid\_loss}$ on the validation set.
C.4 Covariance Parameterization
The covariance matrix $\boldsymbol{\Sigma}_{t}$ is parameterized directly by the forecasting model. Specifically, it is constructed as: $\boldsymbol{\Sigma}_{t}=\boldsymbol{L}_{t}\boldsymbol{L}_{t}^{\top}+\text{diag}(\mathbf{d}_{t}),$ where $\boldsymbol{L}_{t}$ is a low-rank matrix and $\mathbf{d}_{t}$ is a positive diagonal vector. This parameterization ensures that $\boldsymbol{\Sigma}_{t}$ remains positive semi-definite while being computationally efficient to learn. This parameterization is standard in probabilistic forecasting and allows the model to learn both the structure (through $\boldsymbol{L}_{t}$) and scale (through $\mathbf{d}_{t}$) of the covariance during training.
Without constraints, the MVG-CRPS loss could potentially be minimized by driving all eigenvalues of $\boldsymbol{\Sigma}_{t}$ to zero, resulting in a trivial solution. However, this is prevented through the following mechanisms:
  • The diagonal entries of the covariance matrix are parameterized as $d_{i,t}=\mathtt{softplus}(d_{i,t}+\mathtt{diag\_bias})+\sigma_{\text{min}}^{2}$, where the $\mathtt{softplus}$ function ensures that the diagonal entries are strictly positive, regardless of the raw input values, $\mathtt{diag\_bias}$ is initialized to approximately $\mathtt{softplus\_inv}(\sigma_{\text{init}}^{2})$, ensuring that the diagonal entries are initially close to $\sigma_{\text{init}}^{2}$. For instance, with $\sigma_{\text{init}}=1.0$, the initial diagonal values start near 1.0. The addition of $\sigma_{\text{min}}^{2}$ provides a lower bound on the diagonal entries, ensuring that eigenvalues cannot approach zero.
  • The low-rank component is parameterized as $\mathbf{L}_{i,t}=\frac{\mathbf{L}_{i,t}}{\sqrt{R}}$, where dividing by rank ensures that the low-rank term is well-scaled relative to the diagonal entries. This normalization prevents the low-rank component from dominating or becoming disproportionately small in the covariance matrix.
Moreover, the MVG-CRPS loss provides a balance between the calibration and sharpness of the forecasts:
$\mathbf{w}_{t}=\boldsymbol{S}_{t}^{-\frac{1}{2}}\mathbf{v}_{t}=\boldsymbol{S}_{t}^{-\frac{1}{2}}\boldsymbol{U}_{t}^{\top}\left(\mathbf{z}_{t}-\boldsymbol{\mu}_{t}\right),$
$\mathcal{L}=\sum_{t=1}^{T}\sum_{i=1}^{N}\sqrt{\lambda_{t}^{i}}\operatorname{CRPS}\left(\Phi,{w}_{i,t}\right).$
We observe that if the eigenvalues $\lambda_{t}^{i}$ in $\boldsymbol{S}_{t}$ approach zero, $w_{i,t}$ will be scaled very aggressively. This leads to inflated residuals $w_{i,t}$, which subsequently affect the CRPS computation. Since the CRPS metric integrates over the forecast distribution $F(y)$, penalizing deviations between $F(y)$ and the empirical step function $\mathbf{1}(y\geq w_{i,t})$, artificially large $w_{i,t}$ values (resulting from extreme eigenvalue scaling) will cause the CRPS term to increase significantly. This behavior reflects the importance of ensuring that eigenvalues $\lambda_{t}^{i}$ are well-regularized to prevent distortion in the forecast evaluation. By balancing the eigenvalue contributions, the MVG-CRPS ensures both stable calibration and sharpness in probabilistic forecasting.
C.5 SVD and Gradient Calculation
We perform SVD on $\boldsymbol{\Sigma}(\mathbf{h}_{t})$ to obtain $\mathbf{U}_{t}$ and $\mathbf{S}_{t}$ (the eigenvectors and eigenvalues, respectively). These are required to compute the whitening transformation: $\mathbf{w}_{t}=\mathbf{S}_{t}^{-\frac{1}{2}}\mathbf{U}_{t}^{\top}(\mathbf{z}_{t}-\boldsymbol{\mu}_{t}).$ During training, gradients of $\mathcal{L}$ need to flow back through the whitened vecotr $\mathbf{w}_{t}$, the eigenvectors matrix $\mathbf{U}_{t}$, the eigenvalues matrix $\mathbf{S}_{t}$, and the covariance matrix $\boldsymbol{\Sigma}_{t}$. The gradient of $\mathcal{L}$ with respect to $\mathbf{w}_{t}$ is $\frac{\partial\mathcal{L}}{\partial\mathbf{w}_{t}}$. Gradients of $\mathbf{w}_{t}$ are propagated to the whitening transformation: $\mathbf{w}_{t}=\mathbf{S}_{t}^{-\frac{1}{2}}\mathbf{U}_{t}^{\top}(\mathbf{z}_{t}-\boldsymbol{\mu}_{t})$, which involves: (1) gradients with respect to $\mathbf{U}_{t}$; (2) gradients with respect to $\mathbf{S}_{t}^{-\frac{1}{2}}$ (i.e., the square root and inverse of singular values); and (3) gradients with respect to $(\mathbf{z}_{t}-\boldsymbol{\mu}_{t})$. Using PyTorch’s $\mathtt{torch.linalg.svd}$, we calculate the gradients of $\mathbf{U}_{t}$ and $\mathbf{S}_{t}$ via automatic differentiation. For the forward pass, the cost of SVD for $\boldsymbol{\Sigma}(\mathbf{h}_{t})\in\mathbb{R}^{B\times B}$ is $O(B^{3})$, where $B$ is the matrix dimension. For the backward pass, computing the gradients of $\mathbf{U}_{t}$ and $\mathbf{S}_{t}$ also incurs $O(B^{3})$ computational cost. Memory usage scales as $O(B^{2})$ for storing the covariance matrix and the singular value decomposition outputs ($\mathbf{U}_{t},\mathbf{S}_{t}$). Additional memory is required for autograd intermediate values, scaling as $O(B^{3})$. By leveraging PyTorch’s autograd system, we integrate the computation of $\mathbf{U}_{t}$, $\mathbf{S}_{t}$, and their gradients seamlessly into our end-to-end learning pipeline. This ensures that the whitening transformation and the loss function are fully differentiable, allowing the model parameters to be trained via gradient-based optimizers. The parameter $B$ also plays a crucial role in the scalability of our method. By leveraging the Gaussian assumption, we are able to train the model using a much smaller subset of time series at each step. Consequently, the size of the covariance matrix is reduced to $B\times B$, as opposed to $N\times N$, where $N$ represents the total number of time series in the dataset. This design ensures that the computational complexity of our method does not scale with $N$. Moreover, $B$ is kept relatively small in our implementation (e.g., $B=20$), making the approach computationally efficient.
C.6 Naive Baseline Description
In this paper, we use Vector Autoregression (VAR) (Lütkepohl, 2005) as a naive baseline model. The VAR($p$) model is formulated as
$\mathbf{z}_{t}=\mathbf{c}+\boldsymbol{A}_{1}\mathbf{z}_{t-1}+\dots+\boldsymbol{A}_{p}\mathbf{z}_{t-p}+\boldsymbol{\epsilon}_{t},\quad\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\epsilon}),$ (20)
where $A_{i}$ is an $N\times N$ coefficient matrix, and $\mathbf{c}$ is the intercept term. In our experiments, we employ a VAR model with a lag of 1 (VAR(1)). The parameters in Eq. (20) are estimated using ordinary least squares (OLS), as described in Lütkepohl (2005). VAR models are not applied to datasets with insufficient time series in the testing set and are marked as “N/A” in this paper.
Appendix D Additional Results
In Table 6 and Table 7, we compare the energy score performance of models trained with different loss functions.
Table 6: Comparison of energy score across different scoring rules in the multivariate autoregressive forecasting task. The best scores are in boldface. MVG-CRPS scores are underlined when they are not the best overall but exceed the log-score. Mean and standard deviation are reported from 10 runs of each model.
VAR GPVar Transformer
log-score energy score MVG-CRPS log-score energy score MVG-CRPS
$\mathtt{elec\_au}$ (×$\times 10^{3}$) N/A 5.4013$\pm$0.0372 3.9136$\pm$0.0177 4.1508$\pm$0.0283 7.0039$\pm$0.0219 6.3135$\pm$0.0243 3.5217$\pm$0.0150
$\mathtt{cif\_2016}$ (×$\times 10^{3}$) 125.6177$\pm$0.0000 4.2733$\pm$0.0218 4.9329$\pm$0.0161 4.1677$\pm$0.0198 4.6316$\pm$0.0270 4.1063$\pm$0.0241 3.5559$\pm$0.0145
$\mathtt{elec}$ (×$\times 10^{4}$) 10.4788$\pm$0.0757 3.3124$\pm$0.0580 4.8317$\pm$0.0434 3.2435$\pm$0.0300 3.4724$\pm$0.0229 4.3757$\pm$0.0414 3.9672$\pm$0.0374
$\mathtt{elec\_weekly}$ (×$\times 10^{7}$) 2.2191$\pm$0.0308 2.5724$\pm$0.0799 0.8948$\pm$0.02344 1.4040$\pm$0.0887 1.5463$\pm$0.0582 0.9338$\pm$0.0308 0.9985$\pm$0.0360
$\mathtt{exchange\_rate}$ 0.1301$\pm$0.0002 0.3972$\pm$0.0074 0.1895$\pm$0.0034 0.1216$\pm$0.0013 0.2136$\pm$0.0026 0.1774$\pm$0.0026 0.2040$\pm$0.0045
$\mathtt{kdd\_cup}$ (×$\times 10^{2}$) N/A 4.7575$\pm$0.0186 4.3981$\pm$0.0164 4.0719$\pm$0.0180 4.2809$\pm$0.0134 5.9466$\pm$0.0427 3.1788$\pm$0.0122
$\mathtt{m1\_yearly}$ (×$\times 10^{4}$) N/A 7.3860$\pm$0.0789 7.7576$\pm$0.0335 6.1985$\pm$0.0505 8.7079$\pm$0.1760 5.7774$\pm$0.0755 7.5130$\pm$0.1784
$\mathtt{m3\_yearly}$ (×$\times 10^{3}$) N/A 3.6113$\pm$0.0703 2.2147$\pm$0.0427 1.4775$\pm$0.0495 3.1996$\pm$0.0995 4.0982$\pm$0.0621 2.4253$\pm$0.0914
$\mathtt{nn5\_daily}$ (×$\times 10^{2}$) 4.9419$\pm$0.0056 3.3001$\pm$0.0050 3.3004$\pm$0.0052 3.3934$\pm$0.0045 3.2546$\pm$0.0033 3.1622$\pm$0.0045 3.0996$\pm$0.0025
$\mathtt{saugeenday}$ (×$\times 10^{2}$) N/A 1.8098$\pm$0.0231 1.7135$\pm$0.0150 1.9400$\pm$0.0208 1.5780$\pm$0.0183 1.5883$\pm$0.0108 1.8043$\pm$0.0204
$\mathtt{sunspot}$ (×$\times 10$) N/A 2.7737$\pm$0.1195 3.1658$\pm$0.0792 2.6195$\pm$0.1003 5.4893$\pm$0.1132 2.3153$\pm$0.0467 3.2663$\pm$0.0745
$\mathtt{tourism}$ (×$\times 10^{5}$) 3.5958$\pm$0.0354 6.1085$\pm$0.1132 5.6774$\pm$0.0493 5.2111$\pm$0.0896 5.0645$\pm$0.0526 4.7502$\pm$0.0585 5.2702$\pm$0.0853
$\mathtt{traffic\_nips}$ 3358.5004$\pm$10.7535 2.2924$\pm$0.0034 2.1140$\pm$0.0023 2.2916$\pm$0.0015 2.2043$\pm$0.0012 2.2250$\pm$0.0018 2.2000$\pm$0.0018
Table 7: Comparison of energy score across different scoring rules in the univariate Seq2Seq forecasting task.
N-HiTS
log-score energy score MVG-CRPS
$\mathtt{covid}$ (×$\times 10^{5}$) 2.1220$\pm$0.0304 N/A 0.9401$\pm$0.0186
$\mathtt{elec\_hourly}$ (×$\times 10^{5}$) 0.9283$\pm$0.0161 N/A 0.9088$\pm$0.0079
$\mathtt{elec}$ (×$\times 10^{5}$) 0.2535$\pm$0.0018 0.3123$\pm$0.0019 0.2431$\pm$0.0020
$\mathtt{exchange\_rate}$ 0.2876$\pm$0.0055 0.1272$\pm$0.0022 0.1240$\pm$0.0022
$\mathtt{m4\_hourly}$ (×$\times 10^{4}$) 0.2852$\pm$0.0026 0.2890$\pm$0.0029 0.2423$\pm$0.0027
$\mathtt{nn5\_daily}$ (×$\times 10^{3}$) 0.4170$\pm$0.0018 0.3272$\pm$0.0005 0.3958$\pm$0.0021
$\mathtt{pedestrian}$ (×$\times 10^{3}$) 1.1571$\pm$0.0177 0.9746$\pm$0.0081 0.8337$\pm$0.0066
$\mathtt{saugeenday}$ (×$\times 10^{2}$) 1.6690$\pm$0.0391 1.7752$\pm$0.0216 1.7698$\pm$0.0129
$\mathtt{taxi\_30min}$ (×$\times 10^{2}$) 6.9676$\pm$0.0045 6.7906$\pm$0.0058 5.6679$\pm$0.0004
$\mathtt{traffic}$ 3.6810$\pm$0.0136 2.2524$\pm$0.0018 2.2200$\pm$0.0022
$\mathtt{uber\_hourly}$ 6.3252$\pm$0.1785 5.4214$\pm$0.0326 4.2826$\pm$0.0320
$\mathtt{wiki}$ (×$\times 10^{6}$) 1.1535$\pm$0.0047 0.9352$\pm$0.0069 0.9338$\pm$0.0083